OpenStudy (anonymous):

What is the equation of the line that passes through (-2,4) and is perpendicular to 2x+3y=-6

OpenStudy (anonymous):

What are the two pieces of information you need to define a line?

OpenStudy (anonymous):

doesn't say..just says. A.y=-3/2x-2 B.3/2x+7 C.y=-2/3x+8/3 D.y=2/3x-4

OpenStudy (anonymous):

I'm asking in general, for any line, do you know the two pieces of information that you need to define it uniquely?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a point and a slope uniquely identify a line. So if you want to find a line, you need to find a point on the line, and its slope. The problem is asking you to find a line, and you already have a point on it, so you just need to find its slope. does that make sense?

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

well can you find the slope of the line you're given first?

OpenStudy (anonymous):

how would you find the slope of 2x + 3y = -6

OpenStudy (anonymous):

you've already posted a bunch more problems I see. maybe concentrate on one at a time?

OpenStudy (anonymous):

hello?